OpenStudy (anonymous):

Marcel has a data set for which the mean is 33. Each value in the data set is multiplied by 5 what is the mean for the new data set

ganeshie8 (ganeshie8):

consider a dataset : {2, 4}

ganeshie8 (ganeshie8):

its mean = (2 + 4)/2 = 3

ganeshie8 (ganeshie8):

Multiply each element by 5, and see how it affects the mean

ganeshie8 (ganeshie8):

If you will notice, 5 simply factors out from each element

ganeshie8 (ganeshie8):

new dataset : {2*5, 4*5} new mean = (2*5 + 4.5)/2 = 5(2+4)/2 = 5*(old mean)

ganeshie8 (ganeshie8):

In general, multiplying each value in the dataset by some by some number "k" increases the mean by a factor of "k"